# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ Abstract
A “Non-perturbative General Approximation Scheme”(NGAS) is proposed which is potentially applicable to arbitrary interacting quantum systems described by a Hamiltonian. The essential feature of this scheme consists of realization of a “mapping” of the original “interacting system” on to an “exactly solvable” model, while preserving the major effects of interaction through the self consistency requirement of equal quantum averages of observables in the two systems. We apply the method to the different cases of the one dimensional anharmonic-interactions (AHI), which includes the case of the quartic-, sextic- and octic-anharmonic oscillators and quartic-, sextic-double well oscillators within the harmonic approximation and demonstrate how this simple approach reproduces, in the leading order (LO), the results to within a few percent, of some of the earlier methods employing rather different assumptions and often with sophisticated numerical analysis. We demonstrate the flexibility of the proposed scheme by carrying out the analysis of the AHI by choosing the infinite square-well potential (ISWP) in one dimension as the input approximation. We extend the formalism to ${\lambda\phi^{4}}$-quantum field theory (in the massive symmetric-phase) to show the equivalence of the present method to the “Gaussian-effective potential” approach. The structure and stability of the Effective Vacuum is demonstrated by identifying the Boguliubov-Valatin transformation from the’ free theory’ to the ’effective theory’. A prominent feature of NGAS is the emergence of a new formulation of perturbation theory designated as: “Mean Field Perturbation Theory (MFPT)”, which is free from power-series expansion in any physical parameter, including the coupling strength. Its application is thereby extended to deal with interactions of arbitrary strength and to compute system properties having non-analytic dependence on the coupling, thus overcoming the primary limitations of the “standard formulation of perturbation theory” (SFPT). We demonstrate Borel-summability of MFPT for the case of the quartic- and sextic-anharmonic oscillators and the quartic double-well oscillator (QDWO) by obtaining uniformly accurate results for the ground state of the above systems for arbitrary physical values of the coupling strength. The energy-spectrum obtained for the quartic double-well oscillator may be of special significance since “renormalon”-free, unambiguous results are achieved in MFPT, in contrast to the well-known failure of SFPT in this case. The general nature and the simplicity of the formulation of NGAS lead us to conjecture that this scheme may be applicable to arbitrary interactions in quantum theory.
@@
I Introduction : Need for Approximation in Quantum Theory (QT)
It is well known that quantum theory (QT) describes [1,2] the observed features of the physical universe at the fundamental level, both at the microscopic and the macro- scales. However, in contrast to the vastness of the range of application of QT, the exact analytic solutions to the quantum- dynamics of interacting systems are rather sparse. Some of these cases admitting exact solution are described in the texts [3]. It is not difficult to count and classify the cases where exact analytic solutions are possible. In fact, it is believed [4,5] that the existence of exact analytic solutions is related to the factorization property of the corresponding quantum-Hamiltonian. In that context, recent applications of super-symmetric quantum mechanics (SUSYQM) [6] have resulted in considerable extension [6] of the range and class of exactly solvable potentials in non-relativistic QT.
In view of the rather limited range of physical phenomena amenable to exact analytic solutions in QT, it has become inevitable to develop approximation methods of to derive meaningful results. This aspect was realized soon after the discovery of QT [7].
Approximation methods in quantum theory have a long history-the prominent ones being as old as QT itself [3,7]. There is a vast amount of literature already accumulated over the years on the various approximation schemes (AS) proposed. Nevertheless, the list is ever growing with newer proposals. This could be indicative of the fact that an “ideal” approximation scheme (in the sense of general applicability to the multitude of interacting systems in QT) is perhaps yet to be realized.
In order to motivate and understand the scope of the investigations presented in this report, it is therefore, necessary to review and survey the various existing approximation schemes (AS) in QT. This is attempted in the next section
II Survey and classification of the existing approximation schemes
The main approximation - schemes in QT which have become the most useful and found mention in text books [3] are the following: (a)Perturbation theory (b)Variational method and (c) Semi-classical(WKBJ) approximation schemes. There are several other approaches, some of which, are variants of the basic schemes. These are discussed at length in ref.[8]. However, for completeness, we recount these briefly in the following:
@@ II.1 Perturbation theory (PT)
where $g$ is the coupling-strength of interaction, $H_{s}$ is the exactly solvable unperturbed Hamiltonian and $gH_{I}$ is the ( usually non-linear) interaction part of the Hamiltonian. This latter interaction-term prevents exact analytic solution for the eigen-spectrum of $H$. For the eigen-value(s) of $H$, the perturbation expansion, eqn.(1) has the standard interpretation: ${E(g)}$ is the energy of the interacting-system, $\mathcal{E}_{0}$ is the same without interaction, which can be calculated exactly and the $\mathcal{E}_{k}$ represents perturbative- correction at the $k$-th order to the energy $\mathcal{E}_{0}$ of the non-interacting theory. These corrections are computable, in principle, for arbitrary order $k$.
It is clear from the definition, eqn.(1) that perturbative corrections become meaningful only when these remain sub-dominant to the leading order contribution which is the first term on the r.h.s. Furthermore, the coupling strength, $g$ is restricted to small values, i.e. $|g|<1$, in order that the perturbation series does not become a priori meaningless. A further practical requirement is that the sequence represents either convergent or asymptotic series in order that the eqn.(1) makes sense if truncation of the same after a finite number of terms is to yield results approximating the actual value of the observable. An important corollary is, therefore, that the  rate of convergence of the series ( or, the rate of decrease of the sub-asymptotic sequence, as the case may be ) is an important consideration bearing upon the practical application of the method.
Merits and limitations of perturbation theory
Perturbation theory continues to be the preferred method of approximation for various practical and theoretical reasons [9].
Merits
Two main reasons for its popularity could be :
(i)Provision for systematic order-by-order improvement of the leading-order result.
Perturbation theory is defined ( cf. eqn.(1)) as providing systematic corrections order-by-order over a leading order result, which is exactly computable without any approximation. This could be regarded as the main merit of this method.
(ii) The universality of the method
Perturbation Theory is universal in the sense that it is applicable to a general class of interacting systems provided that one can identify an exactly solvable unperturbed system and a perturbed part involving some small parameter for a power-series expansion. This encompasses a wide class of problems in QT including many-body theory and quantum-field theory (QFT).
Limitations of SFPT
The main limitations of the SFPT could be the following:
(i) Inability to describe the so-called “non-perturbative” phenomena
The nature of the power-series expansion (see eqn.(1)) defining SFPT implies that the latter fails to describe a very important class of phenomena when the interaction-strength $g$ is large: $|g|>1$, as happens, e.g. in strong-interactions, bound states etc.
For the same reason, processes/ phenomena involving non-analytic-dependence on $g$ at the origin remain inaccessible to SFPT. There is a large variety of physical phenomena, which fall under this category, e.g. tunneling , decay, phase-transitions, critical-phenomena, collective and cooperative phenomena such as super-conductivity, super-fluidity etc.
The inability to describe the above class of the so-called “non-perturbative” phenomena remains perhaps the most severe limitation of the method.
(ii)The instability of the “perturbative vacuum”
The “perturbative vacuum/ground state” is defined to be the one belonging to the non-interacting theory. In most of the physical processes it turns out that the true (interacting) vacuum/ ground-state is not analytically accessible from the perturbative vacuum/ ground state when interaction is turned on [8,10]. Furthermore, the instability of the perturbative ground-state has been demonstrated in several cases by establishing that the ground state of the interacting theory lies much lower in energy than the former.
(iii) Inconsistency with the known analytic properties
The analytic properties of an observable such as the energy, at the origin in the complex-$g$-plane have been independently established [11] for some systems and these are inconsistent with those implicit in the defining eqn.(1).
(iv) Practical Difficulties
In addition to the above limitations of the SFPT on theoretical grounds, there may be practical difficulties arising, for example, from computing corrections beyond the first few orders, which generally involve the (infinite) sum over the intermediate states. This problem has been addressed by several authors and it has been found that the difficulty can be surmounted , in specific examples , by appealing to special techniques, such as the Dalgrano-Lewis method [12] and the use of the hyper-virial theorem [9].
Other limitations of PT are discussed in subsequent Section(s) in respect of specific systems.
@@
We close this subsection noting that the difficulties encountered with the SFPT have led to several alternative approaches some of which are discussed below.
@@ II.2 Variational Methods (VM)
Thus, the Ritz-method enables one to obtain a least-upper bound (LUB), for a given choice of the trial-state, which is then the closest approximation to the ground-state energy for that choice. Some other relevant aspects of the VMA are discussed below :
(a) The equality sign in eqn.(3) holds if (and only if) the trial-state, $|~{}\psi~{}>~{}$ coincides with the true ground state, $|~{}\psi_{0}~{}>$ of the system.
(b) An error $\sim~{}O(\epsilon)$ in the choice of the trial-state gives rise to an error $\sim~{}O(\epsilon^{2})$ in the computed value of the energy, which means that the system-energy is more accurately determined by the VMA than the wave function.
Consequently, the over-all accuracy of the method sensitively depends on the choice of the trial-state, consistent with the physical boundary conditions for the system. This calls for considerable ingenuity and insight in the choice of the trial-state.
(c) The generalization of the Rayleigh-Ritz method of the VMA to obtain estimates/approximation for the higher excited states can be achieved by several methods. Some of these methods are outlined below :
(i) A tower of trial states, ${}|~{}\psi_{\alpha}~{}>~{};~{}\alpha~{}=~{}1,~{}2,~{}3......,$ approximating the higher excited states of the system may be constructed with the requirement that these are mutually orthogonal to each other, as well as, orthogonal to the variational-ground-state. The construction can be achieved by any suitable method of orthogonalization, such as the Schmidt-method [15] . Denote by $E_{\alpha}$, the energy-functional for the $\alpha$-th state minimized with respect to its free-parameters and then arranged as an decreasing sequence, i.e.,
@@
It is obvious that the above method becomes efficient in practice when the truncation-error (due to the chosen finite value of $k$) becomes small and further that the spectrum stabilizes (with increasing value of $k$). Clearly, therefore, intelligent choice of the basis-states, $|~{}u_{n}>$ becomes crucial in this approach.
@@
The merits and limitations of the VMA may be as described below:
merits
(i) Non-perturbative nature
The method works for arbitrary coupling-strength of interaction as well as for other non-perturbative phenomena with non-analytic dependence on the coupling parameter. This is perhaps the main advantage of VMA.
(ii) Accuracy
Accuracy can be achieved to great precision provided that the trial-states(s) are appropriately chosen. In particular, the diagonalization of the Hamiltonian in a suitably chosen basis-states , with parameters, which are fixed by variational minimization , can attain high accuracy and stability by inclusion of sufficient number of terms in a recursive manner.
(iii) Universality
The method can be applied in principle to arbitrary Hamiltonian systems.
(iv) Establishing vacuum stability
The stability of the ground state/ vacuum of an interacting system can be established by using the basic theorems of the VMA with suitably chosen trial-state. This method has been used e.g. to establish the instability of the perturbative- vacuum in anharmonic-interactions [17], $\lambda\phi^{4}$ QFT [18], QCD [19] etc.
(v)Fail-safe method for guidance
VMA provides guidance to the true ground state and the energy spectrum of the system when no other method is available/ applicable.
@@
(i)Not improvable order-by-order
In the original formulation, this is perhaps the main limitation - unlike PT it is not amenable to systematic improvement order-by-order.
(ii)Non-uniqueness
There is no systematic and standard method to follow in implementation since choice of the trial-states are mainly arbitrary and mostly left to depend upon the ingenuity and insight in the exploration.
(iii) Practical Difficulty in implementation
Practical difficulties may arise in implementation of the VMA when the trial-state(s) involve several variational parameters.
(iv)Not easily generalized to QFT
For QFT, the preferred method of investigation continues to be SFPT mainly because of the advantage of perturbative realization of the renormalization program in a straight forward manner, in spite of valiant attempts within the VMA [18].
There have been several approaches to overcome the limitations of the basic-VMA. These include:(i) improving the input-trial state(s) by either inclusion of more free parameters or otherwise. A sample list can be found in refs.[20-22] for application to standard examples of interacting systems with varying degree of success. However, as noted earlier, these attempts make the variational-minimization of the energy-functional more intractable. (ii)Tightening the variational estimates done by obtaining both an upper-bound as well as, a lower-bound, see e.g. ref.[22]. Lower bounds are obtained by employing standard inequalities and other techniques. (iii)The method of minimum energy variance (MEV): In stead of the Hamiltonian, $H$ itself, upper bounds can be derived, in principle, for any arbitrary function $f(H)$ by the generalization of the standard Rayleigh-Ritz method. When the chosen function is considered as the variance of the Hamiltonian, the method is known as[23] “method of minimum energy variance (MEV)”. The MEV has been shown [23] to lead to improved estimate of the trial-state as well as, the energy eigen-value. In particular, it has been demonstrated[23] that the estimates go beyond the simple Gaussian approximation [17] .
@@ II.3 Semi-Classical Methods (SCM)
where,${}k^{2}(x)~{}=~{}({2m}/{\hbar^{2}})~{}(E~{}-~{}V(x))~{}$; a ,b  are a set of adjacent “turning points” obtained by solving the equation, $k(x)~{}=~{}0~{}$ and other terms have standard meaning. For those cases where the above integral can be evaluated, eqn.(9) can be expressed as :
@@
where, the l.h.s. of eqn.(10) represents the value of the integral as a function of the energy ($E$) and coupling-strengths ${}(\lambda,g....)~{}$ occuring in the potential $V(x)$. The determination of energy $E$ as a function of the other parameters $(n,\lambda,g...)$ therefore, requires the inversion of eqn.(10).
@@ II.4 Other Methods
These are too numerous to be included in this short review. In the following, we succinctly describe the ones that have direct- bearing on the approach presented in this report.
@@
Noting that the VMA and the PT broadly complement each other in terms of advantages and limitations, it is not surprising to find in the literature several attempts to incorporate the essential ingredients of the two methods in a single-framework. We mention some of these in the following discussion.
Several authors have worked using this combined approach, including the following: Halliday and Suranyi [26], W.Caswell [27],J.Killingbeck [28], Hsue and Chern [29], Feynman and Kleinert [30] and collaborators [31], Patnaik [32], Rath [33] and many more. The linear delta-expansion method [34], the Gaussian effective potential (GEP) approach [35], the self-consistent field method [36] and its generalizations [37],methods based upon cluster-expansion [38] , coherent- and squeezed states[39] and several other methods utilize the VMA and the VPA in some form or other. The review of each individual work as mentioned above falls beyond the scope of the present report. However, it may be relevant here to note certain common features of the techniques used by these authors and the consequent achievements attained.
In most of the approaches listed above the basic Hamiltonian of the system is altered by the addition and subtraction of terms involving certain additionalparameters. Perturbation techniques can then be applied with redefinition of the unperturbed Hamiltonian and the perturbation correction. The arbitrary parameters are then fixed order-by-order in the resulting perturbation series, by the variational minimization of energy or by imposing other constraints such as the ”principle of minimum sensitivity”(PMS)[40].The resulting sequence of corrections is often claimed [26,27,32,33] to be convergent. Thus, the problem of convergence of the naive perturbation theory as well as the absence of a built-in mechanism of systematic improvement in the VMA, are overcome to a large extent in these hybrid method, which may be called variation-perturbation method (VPM). Apart from the above common underlying feature, each approach differs in detail, which may be found in the individual reference cited above.
@@
The OMA has been pioneered by Feranchuk and Komarov [41]. The application of the OMA to the anharmonic interaction and to several other problems has been described in ref.[42] , which also provides a guide to earlier works of the authors.
The basic idea of this method is to formulate the problem in the Fock-space of operators, instead of working with the co-ordinates and momenta. In doing so, an arbitrary parameter $\omega$ is introduced into the theory, which is then fixed by standard variational minimization of the energy-functional order-by-order in a modified PT. Details can be found in ref.[41] . The main limitation of the model appears to be the implementation of the scheme e.g. for non-polynomial interactions, interaction representing analytic functions expressible in infinite series, problems in higher dimensions, quantum field theory etc. In such cases, additional assumptions/methods/skills have to be employed [42].
@@
The Path-Integral (PI) formulation of Quantum Theory [66-68] has traditionally served as an important basis to formulate and improve upon approximation methods. It may be recalled that the Feynman-diagram technique, which is perhaps the most popular method of computation in perturbation theory, originated from the PI-approach [69].The approximation methods can be formulated either in the real-time formulation [66,67] or as the imaginary time, Euclidean-formulation [70]. The latter is found to be ideally suited for extending the application to quantum-statistics [69-71]. A variational method was proposed in the Euclidean formalism of the PI in ref.[71] to obtain an approximation to the partition function of the QAHO and thereby, to obtain an estimate of the ground state energy. Application to quantum field theory and statistical physics /critical phenomena are extensively treated in the text [70].In the context of the QAHO-problem, the stationary-phase approximation was used [72] in the PI-formulation to infer the inapplicability of the SFPT due to the existence of an essential singularity at the origin of the coupling strength plane. In ref.[73], variational lower-bound as well as, upper bound on energy of the QAHO have been obtained by the use of inequalities for the partition function. Semi-classical approximation of partition functions for one-dimensional potentials using the PI-formalism has been described in ref.[74].
@@
Various approximation schemes including the “variational perturbation theory”(VPT)[75,78], semi-classical methods [70,78], Stationary-phase approximation [71,78], loop-wise expansion [78], large-order estimate of perturbation theory [70,78], variational estimation using inequalities [71,78] and many more methods [78] have been based upon the PI-formulation. In particular,the multi-instanton effects for potentials with degenerate ground states is described in ref.[76]. In the context of scalar field theory, ref.[77] describes non-standard expansion techniques for the generating functional in the PI formulation. Several other applications in field theory and statistical physics can be found in ref.[70,78].
@@
The main advantages are wide applications [70,78] and non-perturbative nature of the approximation. Major limitation of the PI-formulation include the mathematical complexity involved in evaluating multi-dimensional integrals - even the simple, exactly solvable potentials such as the hydrogen atom problem, require rather complex techniques [78] for implementation. Inclusion of spin, formulation for fermions and for constrained systems pose special problems. In many cases, it is therefore, preferable to use the standard operator based formalism for implementing the approximation methods. Besides, closed-form analytic expressions for the PI can be obtained only for a few problems with interaction when the former can be approximated by a Gaussian about a saddle-point/ stationary-phase point of the classical action and fluctuations thereabout.

### v3

# Towards Nonperturbative Solution of Quantum Dynamics : A Hamiltonian Mean Field Approximation Scheme with Perturbation Theory for Arbitrry Strength of Interaction

@@ Abstract
We introduce a non-perturbative general approximation scheme (NGAS) that can handle interactions of any strength in quantum theory. This approach starts with an input Hamiltonian that can be solved exactly. The interaction effects are then built into this Hamiltonian through a nonlinear feedback enforced by self-consistency conditions. While the method itself is nonperturbative, it can be systematically improved using a new perturbation method called mean field perturbation theory (MFPT) which does not involve power-series expansion in any small parameter. We put this scheme to the test on one-dimensional anharmonic interactions using the harmonic approximation. The results are consistently accurate across various cases including: quartic, sextic, and octic anharmonic oscillators, as well as, the quartic double-well oscillator (QDWO)—even when the coupling strength varies widely. The method’s flexibility is demonstrated when we swap the input Hamiltonian for an infinite-square-well Hamiltonian and still achieve comparable accuracy. When applied to the ${\lambda}{\phi}^{4}$ quantum field theory, this approach aligns with the Gaussian-effective potential method under the harmonic approximation. Beyond that, it reveals the condensate structure of the effective vacuum and highlights the instability of the perturbative ground state. Notably, our ground-state energy results for the QDWO stand in stark contrast to those from standard perturbation theory, where Borel summation fails regardless of coupling strength.
@@
I Introduction : The Role of Approximation in Quantum Theory (QT)
Quantum theory (QT) is widely recognized as a robust framework that elucidates the fundamental behavior of the physical universe, encompassing phenomena from the minutest particles to the expansive scales of the cosmos [1,2]. Despite its extensive applicability, exact analytic solutions for interacting quantum systems remain scarce. While certain special cases have been exactly solved and documented [3], these are exceptions rather than the rule. It is understood [4,5] that exact solvability often hinges on the factorization properties of the corresponding quantum Hamiltonian. In recent years, advances in supersymmetric quantum mechanics (SUSYQM) [6] have significantly broadened the catalog of exactly solvable potentials within non-relativistic QT. Given the limited range of exactly solvable models, approximation methods are indispensable for extracting meaningful insights from quantum systems—a necessity recognized soon after the birth of QT itself [7]. Over decades, numerous approximation schemes have been developed, some tracing back to QT’s earliest days [3,7]. The literature on such methods is vast and continually expanding, reflecting ongoing efforts to discover a broadly applicable, ”ideal” approximation scheme suitable for the diverse array of interacting quantum systems. To set the stage for the investigations detailed in this report, it is crucial to review and contextualize existing approximation approaches. This overview is presented in the following section.
II Overview and Categorization of Existing Approximation Methods
The principal approximation techniques in quantum theory (QT) that have demonstrated significant utility and are frequently cited in textbooks [3] include: (a) Perturbation Theory, (b) Variational Method, and (c) Semi-classical (WKBJ) approximations. Numerous other methods exist, some of which are variations or extensions of these foundational approaches. A comprehensive discussion of these can be found in reference [8]. For completeness, we briefly summarize them here:
@@ II.1 Perturbation theory (PT)
where $g$ is the coupling-strength of interaction, $H_{s}$ is the exactly solvable unperturbed Hamiltonian and $gH_{I}$ is the ( usually non-linear) interaction part of the Hamiltonian. This latter interaction-term prevents exact analytic solution for the eigen-spectrum of $H$. Interpreting the eigenvalues of $H$ through the expansion in equation (1), $E(g)$ is the energy of the full interacting system, $\mathcal{E}_{0}$ is the energy without interaction, and each $\mathcal{E}_{k}$ is the perturbative correction of order $k$ to the unperturbed energy. These corrections can, in principle, be computed to any desired order. For perturbation theory to be meaningful, the corrections must remain smaller than the leading term, which restricts $g$ to small values, typically $|g|<1$. Additionally, the series in equation (1) should either converge or behave asymptotically so that truncating after a finite number of terms provides a good approximation of the true value. The speed of convergence or the decay rate of the asymptotic terms is therefore vital for the method’s practical success.
Strengths and Drawbacks of Perturbation Theory
Perturbation theory is favored for many practical and theoretical reasons [9].
Strengths:
Two key advantages explain its popularity: (i) It allows systematic improvement, providing corrections order-by-order to a leading-order result that can be computed exactly. This capability for controlled refinement is its main strength. (ii) It is broadly applicable, as long as one can identify an exactly solvable unperturbed system and a small coupling parameter suitable for power-series expansion. This universality covers a wide array of quantum problems, including many-body systems and quantum field theory.
Limitations of SFPT:
SFPT is subject to several significant limitations:
(i) Inability to address ”non-perturbative” phenomena: SFPT employs a power-series expansion (refer to eqn.(1)), which becomes ineffective when the interaction strength $g$ is substantial, specifically when $|g|>1$. This scenario frequently arises in contexts involving strong interactions and bound states. SFPT encounters difficulties with processes that exhibit complex dependencies on $g$ from the outset. Numerous physical phenomena, such as tunneling, decay, phase transitions, critical phenomena, and collective behaviors like superconductivity and superfluidity, fall into this category. This represents a substantial challenge for SFPT. (ii) Instability of the ”perturbative vacuum”: The ”perturbative vacuum/ground state” is derived from the non-interacting theory. Typically, the actual (interacting) vacuum/ground state is inaccessible from the perturbative one upon the introduction of interaction [8,10]. Furthermore, the perturbative ground state is unstable due to the significantly lower energy of the ground state in the interacting theory. (iii) Discrepancy with established analytic properties: The analytic properties of an observable, such as energy, in the initial complex-$g$-plane have been determined for certain systems [11], and they do not align with those in eqn.(1). (iv) Practical Challenges: In addition to theoretical issues, SFPT encounters practical challenges, such as computing corrections beyond the initial few orders, which involve an (infinite) summation over intermediate states. This issue has been addressed using specialized methods, such as the Dalgrano-Lewis method [12] and the hyper-virial theorem [9]. Additional limitations of PT are discussed in subsequent sections concerning specific systems.
@@
We conclude this subsection by noting that the challenges associated with SFPT have spurred the development of several alternative approaches, some of which are discussed below.
@@ II.2 Variational Methods (VM)
Thus, the Ritz method enables the determination of a least upper bound (LUB) for a given choice of the trial-state, which represents the closest approximation to the ground-state energy for that choice. Thus, the Ritz method enables the determination of a least upper bound (LUB) for a given choice of the trial-state, which represents the closest approximation to the ground-state energy for that choice.
Further relevant aspects of the VMA are discussed below:
(a) The equality sign in equation (3) is valid if and only if the trial state, $|~\psi~>$, coincides with the true ground state, $|~\psi_{0}~>$, of the system. (b) An error on the order of $O(\epsilon)$ in selecting the trial state results in an error on the order of $O(\epsilon^{2})$ in the computed energy value. This indicates that the system’s energy is determined with greater accuracy by the Variational Method of Approximation (VMA) than the wave function. Consequently, the overall accuracy of the method is highly dependent on the choice of the trial state, which must be consistent with the system’s physical boundary conditions. This necessitates considerable ingenuity and insight in selecting the trial state. (c) The generalization of the Rayleigh-Ritz method of the VMA to estimate or approximate higher excited states can be achieved through several methods. Some of these methods are outlined below: (i) A series of trial states, $|~\psi_{\alpha}~>~;~\alpha~=~1,~2,~3$......, approximating the higher excited states of the system, may be constructed with the requirement that these states are mutually orthogonal to each other and orthogonal to the variational ground state. This construction can be achieved using any suitable orthogonalization method, such as the Schmidt method [15]. Denote by $E_{\alpha}$ the energy functional for the $\alpha$-th state, minimized with respect to its free parameters and then arranged in a decreasing sequence, i.e.,
@@
For practical applications, it is necessary to truncate the determinant at a finite order $k$’. The solution of eqn.(8) typically yields $k$ roots for the energy $E$, which correspond to the approximate energies of the first $k~$’ excited states. It is apparent that this method becomes practically efficient when the truncation error, arising from the selected finite value of $k$, is minimized, and the spectrum stabilizes as $k$ increases. Consequently, the judicious selection of the basis states, $|~u_{n}>$, is essential in this approach.
@@
The merits and limitations of the VMA can be articulated as follows:
Merits (i) Non-perturbative Nature The method is effective for arbitrary coupling strengths of interaction and other non-perturbative phenomena with non-analytic dependence on the coupling parameter. This is arguably the primary advantage of the VMA. (ii) Accuracy High precision can be achieved if the trial states are appropriately selected. Specifically, the diagonalization of the Hamiltonian in a suitably chosen basis, with parameters determined by variational minimization, can achieve high accuracy and stability by including a sufficient number of terms recursively. (iii) Universality The method is applicable, in principle, to arbitrary Hamiltonian systems. (iv) Establishing Vacuum Stability The stability of the ground state or vacuum of an interacting system can be established using the fundamental theorems of the VMA with a suitably chosen trial state. This method has been employed, for instance, to demonstrate the instability of the perturbative vacuum in anharmonic interactions [17], $\lambda\phi^{4}$ QFT [18], QCD [19], etc.
(v)Fail-safe method for guidance The Variational Method Approach (VMA) offers guidance to the true ground state and the energy spectrum of a system when no other method is available or applicable.
@@
(i)Not improvable order-by-order In its original formulation, a primary limitation is that, unlike Perturbation Theory (PT), it does not allow for systematic improvement on an order-by-order basis. (ii)Non-uniqueness There is no systematic or standard method for implementation, as the selection of trial states is largely arbitrary and relies on the ingenuity and insight of the researcher. (iii)Practical Difficulty in implementation Practical challenges may arise in implementing the VMA when trial states involve multiple variational parameters. (iv)Not easily generalized to Quantum Field Theory (QFT) In the context of QFT, the preferred method of investigation remains Standard Field Perturbation Theory (SFPT) due to its straightforward perturbative realization of the renormalization program, despite significant efforts within the VMA [18]. Several approaches have been proposed to address the limitations of the basic VMA. These include: (i) improving the input-trial state(s) by incorporating additional free parameters or other modifications. A sample list can be found in references [20-22] for applications to standard examples of interacting systems with varying degrees of success. However, as previously noted, these attempts render the variational minimization of the energy functional more intractable. (ii)Tightening the variational estimates by obtaining both upper and lower bounds, as seen in reference [22]. Lower bounds are derived using standard inequalities and other techniques. (iii)The method of minimum energy variance (MEV): Instead of the Hamiltonian, $H$, itself, upper bounds can be derived, in principle, for any arbitrary function $f(H)$ through the generalization of the standard Rayleigh-Ritz method. When the chosen function is considered as the variance of the Hamiltonian, the method is referred to as the ”method of minimum energy variance (MEV)” [23]. The MEV has been shown [23] to lead to improved estimates of the trial state as well as the energy eigenvalue. In particular, it has been demonstrated [23] that the estimates surpass the simple Gaussian approximation [17].
@@ II.3 Semi-Classical Methods (SCM)
[…]
@@ II.4 Other Methods
The multitude of methods precludes their exhaustive inclusion in this concise review. Herein, we briefly delineate those methodologies that are directly pertinent to the approach presented in this report.
@@
Given that the Variational Method Approach (VMA) and Perturbation Theory (PT) offer complementary advantages and limitations, it is unsurprising that the literature contains numerous attempts to amalgamate the core elements of both methods into a unified framework. The following discussion highlights some of these efforts. Several scholars have adopted this combined approach, including Halliday and Suranyi [26], W. Caswell [27], J. Killingbeck [28], Hsue and Chern [29], Feynman and Kleinert [30] and collaborators [31], Patnaik [32], Rath [33], among others. The linear delta-expansion method [34], the Gaussian effective potential (GEP) approach [35], the self-consistent field method [36] and its generalizations [37], methods based on cluster-expansion [38], coherent and squeezed states [39], and various other methods incorporate the VMA and the VPA in some capacity. A comprehensive review of each individual work mentioned above is beyond the scope of this report. However, it is pertinent to highlight certain common features of the techniques employed by these authors and the consequent achievements attained.
In most of the approaches listed above, the fundamental Hamiltonian of the system is modified by the addition and subtraction of terms involving certain additional parameters. Perturbation techniques can then be applied with a redefinition of the unperturbed Hamiltonian and the perturbation correction. The arbitrary parameters are subsequently fixed order-by-order in the resulting perturbation series, either by the variational minimization of energy or by imposing other constraints such as the ”principle of minimum sensitivity” (PMS) [40]. The resulting sequence of corrections is often claimed [26,27,32,33] to be convergent. Thus, the problem of convergence in naive perturbation theory, as well as the absence of a built-in mechanism for systematic improvement in the VMA, are largely overcome in these hybrid methods, which may be termed the variation-perturbation method (VPM). Apart from the aforementioned common underlying feature, each approach differs in detail, which may be found in the individual references cited above.
@@
The OMA has been pioneered  by Feranchuk and Komarov [41]. The application of the OMA  to the anharmonic interaction  and to several other problems has been described in ref.[42] , which also provides a guide to earlier works of the authors. The basic idea of this method is to formulate the problem in the Fock-space of operators, instead of working with the co-ordinates and momenta. In doing so, an arbitrary parameter $\omega$ is introduced into the theory, which is then fixed by standard variational minimization of the energy-functional order-by-order in a modified PT. Details can be found in ref.[41] . The main limitation of the model appears to be the implementation of the scheme e.g. for non-polynomial interactions, interaction representing analytic functions expressible in infinite series, problems in higher dimensions, quantum field theory etc. In such cases, additional assumptions/methods/skills have to be employed [42].
@@
The Path-Integral (PI) formulation of Quantum Theory [66-68] has traditionally served as a significant foundation for formulating and refining approximation methods. It is noteworthy that the Feynman-diagram technique, arguably the most popular computational method in perturbation theory, originated from the PI approach [69]. The approximation methods can be formulated either in the real-time formulation [66,67] or as the imaginary time, Euclidean formulation [70]. The latter is particularly well-suited for extending applications to quantum statistics [69-71]. A variational method was proposed in the Euclidean formalism of the PI in ref.[71] to approximate the partition function of the QAHO and thereby estimate the ground state energy. Applications to quantum field theory and statistical physics/critical phenomena are extensively addressed in the text [70]. In the context of the QAHO problem, the stationary-phase approximation was employed [72] in the PI formulation to deduce the inapplicability of the SFPT due to the presence of an essential singularity at the origin of the coupling strength plane. In ref.[73], variational lower and upper bounds on the energy of the QAHO have been obtained using inequalities for the partition function. The semi-classical approximation of partition functions for one-dimensional potentials using the PI formalism has been described in ref.[74].
@@
Various approximation schemes, including ”variational perturbation theory” (VPT) [75,78], semi-classical methods [70,78], stationary-phase approximation [71,78], loop-wise expansion [78], large-order estimates of perturbation theory [70,78], and variational estimation using inequalities [71,78], among others [78], have been developed based on the PI-formulation. Notably, multi-instanton effects for potentials with degenerate ground states are detailed in ref. [76]. In the realm of scalar field theory, ref. [77] discusses non-standard expansion techniques for the generating functional within the PI formulation. Additional applications in field theory and statistical physics are documented in ref. [70,78].
@@
The primary advantages of this approach are its broad applicability [70,78] and its non-perturbative nature. However, a significant limitation of the PI-formulation is the mathematical complexity involved in evaluating multi-dimensional integrals. Even relatively simple, exactly solvable potentials, such as the hydrogen atom problem, necessitate complex techniques [78] for implementation. The inclusion of spin, formulation for fermions, and constrained systems present particular challenges. Consequently, in many instances, it is preferable to employ the standard operator-based formalism for implementing approximation methods. Furthermore, closed-form analytic expressions for the PI can only be derived for a limited number of problems with interaction, where the former can be approximated by a Gaussian around a saddle-point or stationary-phase point of the classical action and its fluctuations.